I would like to know the answer to this problem because I have tried it many times and can't seem to check my answer evenly. Thank you.
The Madison Local High School marching band sold gift wrap to earn money for a band trip to Orlando,Florida .The gift wrap in solid colors sold for $4.00 per roll ,and the print gift wrap sold for $6.00 per roll .The total number of rolls sold was 480 ,and the total amount of money collected was $2340. How many rolls of each of each kind of gift wrap were sold?

Answer by Alan3354(69443)

The Madison Local High School marching band sold gift wrap to earn money for a band trip to Orlando,Florida .The gift wrap in solid colors sold for $4.00 per roll ,and the print gift wrap sold for $6.00 per roll .The total number of rolls sold was 480 ,and the total amount of money collected was $2340. How many rolls of each of each kind of gift wrap were sold?
--------------
s = solid
p = print
-----
s + p = 480
4s + 6p = 2340
---------
s = 480 - p
sub for s
4(480-p) + 6p = 2340
1920 + 2p = 2340
p = 210
s = 270